OpenStudy (anonymous):

Use the quadratic formula to solve the quadratic equation: 2x2 − 3x − 7 = 0

OpenStudy (anonymous):

https://www.youtube.com/user/Hairfunfashion4teens

OpenStudy (anonymous):

since the equation can not be factored, you must use the formula \[x= -b \pm \sqrt{b^2-4ac}\div2a\] a= 2 b = -3, c = -7

OpenStudy (anonymous):

So you problem will look like this: \[x = -(-3) \pm \sqrt{(-3)^2 -4(2)(-7)}\div2(2)\]. Then you simplify: \[3 \pm \sqrt{65}\div 4\]. Your end product would be \[\frac{ 1 }{4}(3\pm \sqrt{65})\]